Write an equation in slope-intercept form for the line that satisfies the following condition.
passes through (9,-9), parallel to the graph of y= 20/3x+10

Answer by ewatrrr(24785)

Hi,
Note: y = mx +b is the slope-intercept form of a line
.
Parallel to y = (20/3)x + 10 tells us the slope of new line Y is also (20/3)
.
Y = (20/3)x + b
.
Pt(9,-9)
substituting these values of x and y in order to find b
-9 = (20/3)*9 + b
.
-9 - 60 = b
b = -69
.
y = (20/3)x - 69